# bibliography.bib

### v1

@@ Abstract
$(\mathfrak{hc}\hat{\zeta})(\wedge^{k}\mathbf{V}_{n})\in\langle\mathfrak{J}_{\lambda^{\prime}}\rangle_{\lambda^{\prime}\in W\omega_{k}}$.
@@
Since $\operatorname{Rep}(\mathbf{G}\times\mathbf{T}^{2})$ act by convolution with objects having a filtration by products of costandard objects ^$\hat{\nabla}_{w}$, Proposition 2.1.1 gives that ${\varpi}_{\textrm{fr}}$ in fact sends $\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}})$ to $\hat{\mathcal{T}}$. Thus, we get a functor
$\varpi_{\textrm{fr}}:\operatorname{Ho}^{b}(\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}}))\to\operatorname{Ho}^{b}(\hat{\mathcal{T}})\simeq\mathcal{H}_{n}.$
Finally, we employ the following proposition from [arkhipovPerverseSheavesAffine2020]. In any tensor category over a field of characteristic zero one can construct a Koszul complex $K_{\phi,d}$ associated to any morphism $\phi:L\to V$ and $d\in\mathbb{Z}^{>0}$, namely the complex
$0\to Sym^{d}(L)\to\dots\to\wedge^{d-i}(V)\otimes Sym^{i}(L)\to\dots\to\wedge^{d-1}(V)\otimes L\to\wedge^{d}(V)\to 0,$
with differential induced by $\phi$.
Let $K_{\lambda}$ be the Koszul complex $K_{w_{\lambda},\dim V_{\lambda}}$ in $\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}})$.
Proposition 4.4.1 ([arkhipovPerverseSheavesAffine2020], Lemma 20).
Assume there is an action of
$a:\operatorname{Ho}^{b}(\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}}))\curvearrowright\mathcal{C}$
on a $\mathbf{K}$-linear idempotent complete triangulated category $\mathcal{C}$, such that $a(K_{\lambda})x=0$ for all $\lambda$ and some object $x\in\operatorname{Ob}(\mathcal{C})$. Then the functor
@@
$F:\mathrm{perf}_{G}({\mathbf{St}_{\mathfrak{g}}})\to\mathcal{C}.$
The functor $\pi_{\textrm{fr}}$ satisfies the condition of the above proposition, because, on an object of the form $\mathbf{V}\otimes\mathcal{O}_{\overline{C}_{\mathbf{St}_{\mathfrak{g}}}},\mathbf{V}\in\operatorname{Rep}(\mathbf{G}\times\mathbf{T}^{2})$, it is given by $\mathbf{V}\otimes\hat{\mathcal{T}}_{w_{0}}$. We have
$\varpi_{\textrm{fr}}(K_{\lambda})=k_{\lambda}\otimes{\mathcal{T}}_{w_{0}}=0,$
where $k_{\lambda}$ is the (obviously acyclic) Koszul complex in $\operatorname{Rep}(\mathbf{G}\times\mathbf{T}^{2})$. Applying Proposition 4.4.1 we get a functor
$\varpi:\mathrm{perf}({\mathbf{St}_{\mathfrak{g}}})\to\mathcal{H}_{n}.$
@@
It is shown in [bezrukavnikovTwoGeometricRealizations2016] that the convolution $\mathcal{F}\star\mathcal{G}$ in $\mathrm{perf}(\mathbf{St}_{\mathfrak{g}})$ can be computed by taking total complex of a bicomplex $\{\mathcal{F}^{i}\star\mathcal{G}^{j}\}$, where $\{\mathcal{F}^{i}\},\{\mathcal{G}^{j}\}$ are complexes representing $\mathcal{F},\mathcal{G}$ in $\operatorname{Ho}^{b}(\mathrm{coh}^{\mathrm{fr}}_{G}({\mathbf{St}_{\mathfrak{g}}}))$. Similarly, the convolution $\mathcal{F}\star\mathcal{G}$ in $\mathcal{H}_{n}$ can be computed by taking the total complex of a bicomplex $\{\mathcal{F}^{i}\star\mathcal{G}^{j}\}$, where $\{\mathcal{F}^{i}\},\{\mathcal{G}^{j}\}$ are complexes representing $\mathcal{F},\mathcal{G}$ in $\operatorname{Ho}^{b}(\hat{\mathcal{T}})$. Thus, we only need to check
$\varpi(\mathcal{O}_{{\mathbf{St}_{\mathfrak{g}}}/G}\star\mathcal{O}_{\mathbf{St}_{\mathfrak{g}}/G})\simeq\hat{\mathcal{T}}_{w_{0}}\star\hat{\mathcal{T}}_{w_{0}},$
which is again a result of loc.cit. ∎
[…]

### v2

@@ Abstract
$\mathfrak{hc}(\hat{\zeta}(\wedge^{k}\mathbf{V}_{n}))\in\langle\mathfrak{J}_{{\boldsymbol{\lambda}}^{\prime}}\rangle_{{\boldsymbol{\lambda}}^{\prime}\in W{\boldsymbol{\omega}}_{k}}$.
@@
Since $\operatorname{Rep}(\mathbf{G}\times\mathbf{T}^{2})$ acts by convolution with objects having a filtration by products of costandard objects $\hat{\nabla}_{w}$, Proposition 2.1.1 gives that ${\varpi}_{\textrm{fr}}$ in fact sends $\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}})$ to $\operatorname{Tilt}^{\wedge}(G)$. Thus, we get a functor
$\varpi_{\textrm{fr}}:\operatorname{Ho}^{b}(\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}}))\to\operatorname{Ho}^{b}({\operatorname{Tilt}^{\wedge}(G)})\simeq\mathcal{H}_{n}.$
Finally, we employ the following proposition from [AB09]. In any tensor category over a field of characteristic zero one can construct a Koszul complex $K_{\phi,d}$ associated with any morphism $\phi\colon L\to V$ and $d\in\mathbb{Z}^{>0}$, namely the complex $0\to Sym^{d}(L)\to\dots\to\wedge^{d-i}(V)\otimes Sym^{i}(L)\to\dots\to\wedge^{d-1}(V)\otimes L\to\wedge^{d}(V)\to 0,$ with differential induced by $\phi$.
Let $K_{{\boldsymbol{\lambda}}}$ be the Koszul complex $K_{w_{{\boldsymbol{\lambda}}},\dim\bf{V}_{{\boldsymbol{\lambda}}}}$ in $\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}})$. We will need the following proposition, [AB09, Lemma 20].
Proposition 4.4.1.
Assume there is an action
$a\colon\operatorname{Ho}^{b}(\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}}))\to\mathrm{End}(\mathcal{C})$
of $\operatorname{Ho}^{b}(\mathrm{coh}^{\mathrm{fr}}_{\mathbf{G}\times\mathbf{T}^{2}}(\overline{\mathbf{C}}_{\mathbf{St}_{\mathfrak{g}}}))$ on a $\mathbf{K}$-linear idempotent-complete triangulated category $\mathcal{C}$, such that $a(K_{{\boldsymbol{\lambda}}})x=0$ for all $\boldsymbol{\lambda}$} and some object $x\in\operatorname{Ob}(\mathcal{C})$. Then the functor
@@
$F:\mathrm{perf}_{\mathbf{G}}({\mathbf{St}_{\mathfrak{g}}})\to\mathcal{C}.$
The functor $\varpi_{\textrm{fr}}$ satisfies the condition of the above proposition, because, on an object of the form $\mathbf{W}\otimes\mathcal{O}_{\overline{C}_{\mathbf{St}_{\mathfrak{g}}}}$ with $\mathbf{W}\in\operatorname{Rep}(\mathbf{G}\times\mathbf{T}^{2})$, it is given by $\mathbf{W}\otimes\hat{\mathcal{T}}_{w_{0}}$. We have
$\varpi_{\textrm{fr}}(K_{{\boldsymbol{\lambda}}})=k_{{\boldsymbol{\lambda}}}\otimes{\mathcal{T}}_{w_{0}}=0,$
where $k_{{\boldsymbol{\lambda}}}$ is the (obviously acyclic) Koszul complex in $\operatorname{Rep}(\mathbf{G}\times\mathbf{T}^{2})$. Applying Proposition 4.4.1 we get a functor
$\varpi:\mathrm{perf}_{\mathbf{G}}({\mathbf{St}_{\mathfrak{g}}})\to\mathcal{H}_{n}.$
@@
We can identify the category $\mathcal{H}_{n}^{\mathrm{perf}}$ with the bounded homotopy category of free $\mathcal{O}_{\hat{0}}(\mathfrak{t}\times_{\mathfrak{t}/W}{\mathfrak{t}})$-modules, where the subscript in $\mathcal{O}_{\hat{0}}$ stands for the formal completion at $(0,0)\in\mathfrak{t}\times_{\mathfrak{t}/W}{\mathfrak{t}}$ [BR22, Theorem 9.1]. From the computation above it is easy to see that, passing through this equivalence, $\varpi$ can be identified with the functor of restriction to the Steinberg cross-section $\mathfrak{t}\times_{\mathfrak{t}/W}\mathfrak{t}\to\mathbf{St}_{\mathfrak{g}}$. Since the latter is monoidal, and the image of $\mathrm{perf}_{\mathbf{G}}(\mathbf{St}_{\mathfrak{g}})$ is in $\mathcal{H}_{n}^{\mathrm{perf}}$, we get the result.